OpenStudy (anonymous):

Which statement best describes the graph of x3 - 3x2 - 4x + 12?

OpenStudy (anonymous):

It starts down on the left and goes up on the right and intersects the x-axis at x = -2, 2, and 3. It starts down on the left and goes up on the right and intersects the axis at x = -2, 4, and 3. It starts up on the left and goes down on the right and intersects the axis at x = -2, 2, and 3. It starts up on the left and goes down on the right and intersects the axis at x = -2, 4, and 3.

OpenStudy (anonymous):

@ninjasandtigers

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks one more please If (x + 3)2 = 42, what is the approximate positive value of x? 0.48 3.48 5.48 6.48

OpenStudy (anonymous):

(x+3)(x+3)=42 x^2+6x+9=42 x^2+6x-33=0 use the quadratic formula

OpenStudy (anonymous):

okkk thanks

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

oh my gosh, this may sound weird but you remind me of my friend because of your username :p @liorafanny

OpenStudy (anonymous):

hahaha thats funny

OpenStudy (anonymous):

btw i still cant figure out the answer :/

OpenStudy (anonymous):

3.48

OpenStudy (anonymous):

ok and another question

OpenStudy (anonymous):

A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t2 + 32t + 90 The average rate of change of f(t) from t = 4 seconds to t = 6 seconds is _____ feet per second.

OpenStudy (anonymous):

sorry but I really have to take my test and I'm being timed >_<

OpenStudy (anonymous):

kk its fine